OpenStudy (anonymous):

Nora’s leaf collection consisted of 150 leaves in August. By December, her collection had grown to 177 leaves. By how much did her collection increase between August and December? A. 9% B. 12% C. 15% D. 18%

OpenStudy (***[isuru]***):

the amount increased = 177- 150 = 27 the initial amount = 150 \[percentage \ increase = \frac{ the \ amount \ increased }{ the \ initial \ amount}\times 100 = \frac{ 27 }{150 }\times 100\]

OpenStudy (***[isuru]***):

which is 18% :)

OpenStudy (***[isuru]***):

got it :)